Considering the relationships between CT, alkalinity, PCO2, and pH and assuming near-neutral pH and no solids.

a. A groundwater with an alkalinity of 200 mg/L as CaCO3 is in equilibrium with 0.02 atm of CO2. Find the pH and CT of the water.

b. If the groundwater in part a is aerated by equilibration with the atmosphere, what is the resulting pH and CT.
Answer #1

(a): Given the pressure of CO2, P = 0.02 atm

Hecne according to Henry's law,

concentration of CO2 in water, [CO2] = [H2CO3] = KH x P = 0.034 mol.L-1.atm-1x0.02 atm = 6.8 x 10-4 M

Now for the 1st dissociation of H2CO3

For the second dissociation of H2CO3

-S Initial conc(M): 1.IAylo ๕49メ104Hence pH of water is 4.77

Given alkalinity of water is 200 mg/L as CaCO3 = 0.2 g/L

Molecular mass of CaCO3 = 100 g/mol

Hence concentration of CO32-, in water [CO32-] = 0.2 g / 100g/mol / 1L = 2x10-3 M

Hence total  [CO32-]t = 2x10-3 M + 4.8 x 10-11 M \approx  2x10-3 M

Hence CT = [H2CO3] + [HCO3-] + [CO32-]t =  6.63 x10-4 M +  1.69x10-5 M + 2x10-3 M

=  2.647 x10-3 M (answer)

(b): For normal atmosphere, the partial pressure of CO2 is P = 3.95 x10-4 atm

Hecne according to Henry's law,

concentration of CO2 in water, [CO2] = [H2CO3] = KH x P = 0.034 mol.L-1.atm-1x3.95 x10-4 atm = 1.343 x 10-5 M

=> [H2CO3] = 1.343 x 10-5 M

Now use the above concentration of H2CO3 to calculated the pH and CT in a similar way as we did for (a)